A company is trying to expose as many people as possible to its new brand through social media advertising targeting 4 million possible viewers. A model for the number of people P, in millions, who are aware of the new brand after t days of advertising is estimated to be ? = 4(1 – ?−0.037?) a. How many persons can they expect to meet in 1 week? b. How many days, to the nearest day, will the advertising campaign be launched that 60% of the possible viewers will be aware of the new brand?
Answer #1

iwem quation ia, Thu P = 4(1-e0**) where Pin mumber of people im millions and a - o-o 37x7 P()4-e ) Now = 4 (-077123 023) 0.O 037T T 6.0*7 2 T= 24 76 (approximately) T - 25 d ays the Atter o the poible 25 e tt mew brand, aware